OpenStudy (vera_ewing):

Which of the following half-reactions would occur at the anode of an electrochemical cell? A. H+(aq) + Cl− (aq) → HCl(aq) B. Cu(s) → Cu+(aq) + e− C. Cu+(aq) + e− → Cu(s) D. Li+(aq) + e− → Li(s)

OpenStudy (vera_ewing):

@aaronq

OpenStudy (vera_ewing):

@Jesstho.-. help!!!

OpenStudy (aaronq):

so (from the previous question) what happens at the anode?

OpenStudy (vera_ewing):

oxidation

OpenStudy (aaronq):

right, so which one of these equations is an oxidation reaction?

OpenStudy (vera_ewing):

B?

OpenStudy (anonymous):

Exactly. The half-reaction on the anode, where oxidation occurs, is Zn(s) = Zn2+(aq) + 2e-. The zinc loses two electrons to form Zn2+. The half-reaction on the cathode where reduction occurs is Cu2+(aq) + 2e- = Cu(s).

OpenStudy (aaronq):

yup!

OpenStudy (vera_ewing):

so the answer is B?

OpenStudy (anonymous):

The entire reaction can be written by combining both half-reactions: Zn(s) + Cu2+(aq) = Zn2+(aq) + Cu(s)

OpenStudy (anonymous):

Understand?

OpenStudy (vera_ewing):

yep! the answer is B! thank you! :)

OpenStudy (anonymous):

No problem